# VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents

### v1

@@ Abstract
Recent advances in large audio language models (LALMs) have greatly enhanced multimodal conversational systems. However, existing benchmarks remain limited—they are mainly English-centric, rely on synthetic speech, and lack comprehensive, discriminative evaluation across multiple dimensions. To address these gaps, we present Voice Chat Bot Bench (VCB Bench)—a high-quality Chinese benchmark built entirely on real human speech. VCB Bench evaluates LALMs from three complementary perspectives: instruction following (including speech-level control beyond text commands), knowledge understanding (general knowledge, reasoning, and daily dialogue), and robustness (stability under perturbations in content, environment, and speaker traits). Experiments on representative LALMs reveal notable performance gaps and highlight future directions for improvement. VCB Bench provides a reproducible and fine-grained evaluation framework, offering standardized methodology and practical insights for advancing Chinese voice conversational models.  [^1] Code and data are available at https://github.com/193746/VCB-Bench-Evalkit
@@ 3.2 Dataset Details
Model Instruction Following Konwledge
TIF TIF-En SIF SIF-En MTD GK ML DC
GLM4-Voice (Zeng et al., 2024) 85.82 82.52 85.57 78.52 85.13 45.53 62.14 48.64
Kimi-Audio (Ding et al., 2025) 85.13 88.92 85.69 61.87 85.67 53.51 79.94 74.76
Qwen2.5-Omni (Xu et al., 2025a) 87.40 72.58 71.37 58.09 86.93 55.43 80.24 73.72
Baichuan-Audio-Chat (Li et al., 2025) 72.49 76.22 78.96 68.07 73.27 44.48 60.33 54.38
Qwen2-Audio-Instruct (Chu et al., 2024) 84.56 75.86 / / 85.67 35.83 60.78 67.07
StepAudio (Huang et al., 2025) 87.17 66.92 80.25 63.63 / 60.42 77.07 59.52
StepAudio2Mini (Wu et al., 2025) 82.79 75.54 78.81 65.15 87.80 61.15 81.30 83.08
Mimo-Audio 91.21 91.76 72.89 24.25 / 56.58 84.01 87.92
GPT4o-Audio 91.24 91.66 88.15 86.07 / 61.29 77.68 77.64
@@
CV introduce linguistic disruptions: (1) Fillers: incorporate discourse markers (e.g., "um", "ah"); (2) Repetition: include repeated phrases/words; (3) Mispronunciation: introduce phonetic deviations; (4) Grammatical Error: employ ungrammatical constructions; (5) Topic Shift: implement abrupt topic changes; (6) Code-Switching: mix Chinese and English. Each category evaluates the model’s ability to maintain comprehension despite content imperfections.
@@ 4.1 Configuration
We evaluate the latest and most capable LALMs. The selected models comprise GLM4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio-Chat, Qwen2-Audio-Instruct, StepAudio, StepAudio2Mini, Mimo-Audio, and GPT4o-Audio.
@@ 4.2 Main Result
As shown in Table 1, GPT-4o-Audio, as a non-open-source state-of-the-art (SOTA) model, serves as a strong baseline and demonstrates all-round superiority across most tasks, which is reasonable given its non-open nature and advanced proprietary capabilities. Focusing on other open-source E2E LALMs, notable performance discrepancies emerge:
@@
In Instruction Following, Mimo-Audio excels in TIF/TIF-En with scores close to GPT-4o-Audio, indicating strong cross-lingual text instruction adaptation. For SIF, StepAudio series and Kimi-Audio perform robustly in Chinese SIF, yet their SIF-En scores lag significantly behind, reflecting challenges in handling English speech’s paralinguistic features. In MTD, StepAudio2Mini leads among open-source models with 87.80, outperforming counterparts like Baichuan-Audio-Chat, which highlights divergence in long-context dialogue logic control.
In Knowledge, Mimo-Audio stands out in ML (84.01) and DC (87.92), surpassing other open-source models and even GPT-4o-Audio’s performance—suggesting strengths in deep reasoning and text semantic analysis. However, models like Baichuan-Audio-Chat show limited performance in GK (44.48) and ML (60.33), revealing gaps in multi-disciplinary knowledge coverage and step-by-step reasoning.
Overall, open-source E2E LALMs exhibit task-specific strengths (e.g., Mimo-Audio in reasoning, StepAudio2Mini in Chinese dialogue and general knowledge) but face challenges in cross-lingual speech adaptation and comprehensive knowledge reasoning.
@@ 4.4 Pretraining Evaluation
Baichuan-Audio-Base S -> T 52.36 54.41 32.65 58.33
S -> S 25.39 20.69 40.82 31.94
Kimi-Audio-Base S -> T 78.01 76.25 73.47 87.50
S -> S 54.71 49.42 69.39 63.89
Qwen2-Audio-Base S -> T 48.95 51.72 30.61 51.39
S -> S 36.91 36.78 44.90 31.94
StepAudio2Mini-Base S -> T 50.26 52.87 26.53 56.94
S -> S 30.63 27.55 34.69 38.89
@@
The SC task evaluates pre-trained LALMs’ “intelligence” and cross-modal semantic coherence by judging the rationality of story endings. The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in S->T and 54.71 in S->S, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios. In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base score much lower. Moreover, all models perform worse in S->S than S->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
@@ A.2 Complete Results Of Instruction Following
A.3 Complete Results Of Konwledge
Table 14-16 shows the complete results of konwledge. For General Knowledge, GPT4o-Audio (61.29) and StepAudio2Mini (61.15) achieve relatively high average scores. For example, GPT4o-Audio excels in Econ (85.42) and Geogr (62.00), while StepAuido2Mini leads in Chem (80.43) and Phys (74.51). In contrast, Baichuan-Audio-Chat scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
@@ A.5 Complete Results Of Text-Speech Alignment
Qwen2-Audio-Instruct 75.86 66.20 76.20 69.40 79.00 71.80 88.60 79.00
StepAudio 66.92 71.60 72.00 51.60 68.00 69.80 50.60 85.60
@@
GLM4-Voice 78.52 82.80 68.00 64.40 81.40 84.60 88.20
Kimi-Audio 61.87 69.20 54.00 50.00 61.40 67.40 68.40
@@
StepAudio 63.63 64.80 54.20 46.20 75.60 62.80 71.40
StepAudio2Mini 65.15 73.00 47.20 52.40 78.80 70.20 68.00
@@
GLM4-Voice 85.13 92.20 87.20 76.00
Kimi-Audio 85.67 92.40 89.60 75.00
Qwen2-Audio-Instruct 85.67 93.20 91.00 72.80
Qwen2.5-Omni 86.93 93.60 88.80 78.40
Baichuan-Audio-Chat 73.27 80.00 74.60 65.20
StepAudio2Mini 87.80 92.60 94.40 76.20
@@
Qwen2-Audio-Instruct 60.78 86.30 57.06 39.29 51.56 22.50 59.75
StepAudio 77.07 91.10 88.82 75.00 65.63 32.50 68.55
StepAudio2Mini 81.30 97.26 88.82 65.48 71.88 45.00 79.87
Mimo-Audio 84.01 94.52 88.82 72.62 78.13 52.50 85.53
@@
GLM4-Voice 65.00 (-17.20) 83.00 (-1.20) 47.60 (-32.40) 82.40 (0.20) 85.00 (-0.60) 82.60 (-2.60) 83.20 (3.20) 79.40 (0.00) 96.40 (0.60) 96.80 (0.80) 96.00 (0.40) 39.20 (-55.20)
Kimi-Audio 45.60 (-30.80) 66.00 (-7.20) 25.80 (-53.80) 56.20 (-20.60) 54.40 (-23.80) 77.40 (18.80) 73.60 (-7.20) 36.60 (-44.00) 93.80 (-1.60) 93.20 (-2.00) 94.40 (-1.20) 87.80 (-4.40)
Qwen2.5-Omni 75.60 (-5.40) 82.60 (1.40) 68.80 (-12.00) 77.80 (-3.40) 79.40 (-1.80) 81.40 (5.40) 74.40 (-8.00) 77.20 (-5.60) 90.40 (0.20) 93.20 (3.20) 87.60 (-2.80) 89.20 (-1.60)
Baichuan-Audio-Chat 60.40 (-0.80) 61.80 (-0.40) 59.20 (-1.20) 61.60 (-0.60) 61.80 (-0.80) 62.60 (-5.40) 68.00 (3.20) 56.60 (-1.00) 77.20 (3.00) 83.60 (13.20) 70.80 (-7.20) 60.00 (-7.80)
Qwen2-Audio-Instruct 76.40 (-5.80) 80.40 (-1.40) 72.60 (-10.00) 81.60 (0.80) 79.40 (0.60) 80.00 (-6.60) 83.20 (0.00) 83.40 (5.00) 93.20 (-1.40) 93.60 (-1.20) 92.80 (-1.60) 77.80 (-7.00)
StepAudio2Mini 75.60 (-4.80) 75.40 (-2.00) 75.80 (-7.60) 75.40 (-2.00) 78.80 (-0.60) 88.00 (2.60) 67.20 (-4.80) 72.80 (-3.40) 96.60 (2.80) 96.00 (2.00) 97.20 (3.60) 76.00 (-14.40)
Mimo-Audio 76.00 (-14.80) 92.80 (-0.40) 59.60 (-28.80) 88.40 (-0.20) 90.60 (2.40) 96.00 (10.60) 88.80 (1.60) 82.80 (-8.40) 99.20 (1.00) 98.80 (-0.40) 99.60 (2.40) 60.40 (-33.00)
GPT4o-Audio 79.00 (-8.60) 89.00 (0.00) 69.20 (-17.00) 88.40 (2.20) 91.80 (4.20) 89.40 (0.00) 85.60 (0.80) 86.60 (2.20) 96.80 (-1.20) 97.20 (0.00) 96.40 (-2.40) 43.40 (-48.40)
@@
GLM4-Voice 64.00 (-24.60) 45.00 (-36.80) 96.60 (-0.60) 74.80 (-12.60) 89.20 (-0.80) 85.60 (-3.80) 83.00 (-1.40) 93.00 (-0.80) 92.80 (1.60) 78.40 (-14.60)
Kimi-Audio 74.00 (-5.40) 60.60 (-11.60) 94.80 (4.60) 84.00 (-1.20) 75.80 (-9.60) 76.80 (-8.20) 58.20 (-11.60) 93.80 (0.60) 70.20 (-19.80) 76.80 (-11.20)
Qwen2.5-Omni 69.40 (-16.80) 52.00 (-30.60) 95.40 (0.60) 82.80 (-3.00) 85.20 (-4.80) 80.80 (-6.20) 74.80 (-8.20) 95.60 (-1.80) 85.60 (-4.40) 79.80 (-10.20)
Baichuan-Audio-Chat 56.60 (-15.60) 45.00 (-20.00) 81.20 (-8.60) 58.20 (-12.60) 73.40 (-2.80) 73.40 (0.40) 64.40 (1.00) 81.40 (-5.40) 72.00 (-5.60) 67.20 (-4.60)
Qwen2-Audio-Instruct 59.80 (-17.60) 37.80 (-32.00) 93.80 (-1.60) 76.00 (-1.20) 87.60 (-1.60) 82.20 (-6.00) 80.00 (-1.40) 96.20 (-0.20) 88.60 (0.40) 77.80 (-13.40)
StepAudio2Mini 65.00 (-19.20) 45.80 (-33.00) 94.80 (-1.20) 78.80 (-6.40) 83.40 (-5.80) 84.00 (-2.20) 66.20 (-12.00) 94.80 (-2.60) 84.80 (-6.80) 81.60 (-8.40)
Mimo-Audio 73.20 (-20.20) 54.00 (-35.60) 99.40 (-0.60) 90.80 (-4.60) 90.40 (-5.00) 87.00 (-9.00) 86.80 (-4.80) 99.60 (0.60) 85.60 (-8.20) 95.20 (-4.20)
GPT4o-Audio 68.20 (-19.20) 61.00 (-20.60) 98.20 (-0.60) 54.80 (-34.40) 92.00 (-1.60) 86.20 (-6.60) 88.00 (-0.20) 96.40 (0.20) 94.20 (-1.60) 74.60 (-21.60)
@@
Fillers Repetition Mispronunciation Grammatical Error Topic Shift Code Switching
GLM4-Voice 81.80 (-3.00) 89.20 (-0.60) 82.20 (-10.60) 85.60 (-6.00) 86.20 (-6.40) 80.80 (-10.80)
Kimi-Audio 76.00 (-0.40) 79.40 (-1.20) 61.80 (-23.20) 77.40 (-10.40) 79.40 (-6.80) 84.00 (-8.80)
Qwen2.5-Omni 79.00 (-2.00) 84.80 (0.00) 80.60 (-8.40) 79.20 (-8.60) 75.00 (-11.40) 80.40 (-9.20)
Baichuan-Audio-Chat 69.60 (1.20) 72.00 (-0.20) 71.40 (-1.20) 65.60 (-11.00) 61.60 (-12.40) 68.40 (-7.60)
Qwen2-Audio-Instruct 75.00 (1.60) 85.80 (-1.20) 78.40 (-10.00) 79.40 (-9.20) 67.00 (-19.80) 76.40 (-11.60)
StepAudio2Mini 79.20 (-3.80) 81.60 (-2.60) 82.20 (-5.60) 80.80 (-8.20) 80.60 (-3.00) 83.60 (-8.60)
Mimo-Audio 92.00 (-1.80) 95.00 (-0.40) 96.00 (-3.60) 97.40 (-0.60) 92.00 (-5.60) 92.60 (-7.00)
GPT4o-Audio 88.40 (1.80) 93.20 (0.00) 92.60 (-4.00) 89.00 (-5.00) 90.60 (-2.60) 96.80 (-0.60)

### v2

@@ Abstract
Recent advances in large audio language models (LALMs) have greatly enhanced multimodal conversational systems. However, existing benchmarks remain limited—they are mainly English-centric, rely on synthetic speech, and lack comprehensive, discriminative evaluation across multiple dimensions. To address these gaps, we present Voice Chat Bot Bench (VCB Bench)—a high-quality Chinese benchmark built entirely on real human speech. VCB Bench evaluates LALMs from three complementary perspectives: instruction following (including speech-level control beyond text commands), knowledge understanding (general knowledge, reasoning, and daily dialogue), and robustness (stability under perturbations in content, environment, and speaker traits). Experiments on representative LALMs reveal notable performance gaps and highlight future directions for improvement. VCB Bench provides a reproducible and fine-grained evaluation framework, offering standardized methodology and practical insights for advancing Chinese voice conversational models.  [^1] Code and data are available at https://github.com/Tencent/VCB-Bench
@@ 3.2 Dataset Details
Model Instruction Following Knowledge Robustness
TIF TIF-En SIF SIF-En MTD GK ML DC SV EV CV
GLM4-Voice (Zeng et al., 2024) 85.82 82.52 85.57 78.53 85.05 45.53 62.14 48.64 76.09 80.25 84.41
Kimi-Audio (Ding et al., 2025) 85.13 88.92 85.69 61.86 85.66 53.51 79.94 74.76 68.25 75.44 77.71
Qwen2.5-Omni (Xu et al., 2025a) 87.40 72.58 71.37 58.09 86.85 55.43 80.39 73.72 82.31 79.81 80.44
Baichuan-Audio-Chat (Li et al., 2025) 72.49 76.22 78.96 68.07 73.20 44.48 60.33 54.38 65.53 67.62 67.46
Qwen2-Audio-Instruct (Chu et al., 2024) 84.56 75.89 / / 85.57 35.83 59.28 67.07 83.01 78.11 79.60
StepAudio (Huang et al., 2025) 87.16 66.93 80.25 63.66 / 60.42 77.08 59.52 71.63 74.87 85.81
StepAudio2Mini (Wu et al., 2025) 82.79 75.54 78.81 65.14 87.72 61.15 81.29 83.08 81.61 77.84 83.04
Mimo-Audio 91.22 91.76 72.89 24.25 / 56.58 84.01 87.92 84.43 86.19 95.10
GPT4o-Audio 91.24 91.66 88.16 86.07 / 61.29 77.68 77.64 82.32 82.49 92.34
Qwen3-Omni (Xu et al., 2025b) 95.59 91.24 84.37 77.50 91.13 74.74 94.27 94.26 93.58 92.68 97.61
@@
CV introduce linguistic disruptions: (1) Casual Talk: incorporate discourse markers (e.g., "um", "ah") and repeated phrases/words; (2) Mispronunciation: introduce phonetic deviations; (3) Grammatical Error: employ ungrammatical constructions; (4) Topic Shift: implement abrupt topic changes; (5) Code-Switching: mix Chinese and English. Each category evaluates the model’s ability to maintain comprehension despite content imperfections.
@@ 4.1 Configuration
We evaluate the latest and most capable LALMs. The selected models comprise GLM4-Voice, Kimi-Audio, Qwen2.5-Omni, Baichuan-Audio-Chat, Qwen2-Audio-Instruct, StepAudio, StepAudio2Mini, Mimo-Audio, GPT4o-Audio and Qwen3-Omni [^2] Qwen3-Omni was included solely in the experimental comparison in Section 4.2..
@@ 4.2 Main Result
Table 1 presents the average scores of each model across various test sets in VCB Bench, with the detailed scores for all subsets provided in the appendix. As shown, Qwen3-Omni emerges as a new state-of-the-art (SOTA) model, demonstrating all-round superiority across most tasks. GPT-4o-Audio, as a strong non-open-source model, also performs robustly, particularly in SIF.
@@
In Instruction Following, Mimo-Audio excels in TIF (91.22) and TIF-En (91.76) with scores close to GPT-4o-Audio, indicating strong cross-lingual text instruction adaptation. For SIF, GPT-4o-Audio leads in Chinese SIF (88.16) and SIF-En (86.07), while StepAudio series and Kimi-Audio perform robustly in Chinese SIF but their SIF-En scores lag significantly, reflecting challenges in handling English speech’s paralinguistic features. In MTD, StepAudio2Mini (87.72) and Qwen3-Omni (91.13) lead among models, outperforming counterparts like Baichuan-Audio-Chat (73.20), which highlights divergence in long-context dialogue logic control.
In Knowledge, Qwen3-Omni stands out in GK (74.74), ML (94.27), and DC (94.26), surpassing other models and demonstrating strengths in multi-disciplinary knowledge and deep reasoning. Mimo-Audio also performs strongly in ML (84.01) and DC (87.92), suggesting capabilities in deep reasoning and text semantic analysis. However, models like Baichuan-Audio-Chat show limited performance in GK (44.48) and ML (60.33), revealing gaps in multi-disciplinary knowledge coverage and step-by-step reasoning.
Overall, Qwen3-Omni demonstrates comprehensive superiority, while other open-source E2E LALMs exhibit task-specific strengths (e.g., Mimo-Audio in reasoning, StepAudio2Mini in Chinese dialogue and general knowledge) but face challenges in cross-lingual speech adaptation.
@@ 4.4 Pretraining Evaluation
Baichuan-Audio-Base A -> T 52.36 54.41 32.65 58.33
A -> A 25.39 20.69 40.82 31.94
Kimi-Audio-Base A -> T 78.01 76.25 73.47 87.50
A -> A 54.71 49.42 69.39 63.89
Qwen2-Audio-Base A -> T 48.95 51.72 30.61 51.39
A -> A 36.91 36.78 44.90 31.94
StepAudio2Mini-Base A -> T 50.26 52.87 26.53 56.94
A -> A 30.63 27.55 34.69 38.89
@@
The SC task evaluates pre-trained LALMs’ intelligence and cross-modal semantic coherence by judging the rationality of story endings. The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in A->T and 54.71 in A->A, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios. In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base score much lower. Moreover, all models perform worse in A->A than A->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
@@ A.2 Complete Results Of Instruction Following
A.3 Complete Results Of Knowledge
Table 14-16 shows the complete results of knowledge. For General Knowledge, GPT4o-Audio (61.29) and StepAudio2Mini (61.15) achieve relatively high average scores. For example, GPT4o-Audio excels in Econ (85.42) and Geogr (62.00), while StepAuido2Mini leads in Chem (80.43) and Phys (74.51). In contrast, Baichuan-Audio-Chat scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
@@ A.5 Complete Results Of Text-Speech Alignment
Qwen2-Audio-Instruct 75.89 66.20 76.20 69.40 79.00 71.80 88.60 79.00
StepAudio 66.93 71.60 72.00 51.60 68.00 69.80 50.60 85.60
@@
GLM4-Voice 78.53 82.80 68.00 64.40 81.40 84.60 88.20
Kimi-Audio 61.86 69.20 54.00 50.00 61.40 67.40 68.40
@@
StepAudio 63.66 64.80 54.20 46.20 75.60 62.80 71.40
StepAudio2Mini 65.14 73.00 47.20 52.40 78.80 70.20 68.00
@@
GLM4-Voice 85.05 92.12 87.17 76.04
Kimi-Audio 85.66 92.38 89.53 75.06
Qwen2-Audio-Instruct 85.57 93.25 90,91 72.82
Qwen2.5-Omni 86.85 93.69 88.81 78.39
Baichuan-Audio-Chat 73.20 79.91 74.64 65.24
StepAudio2Mini 87.72 92.58 94.44 76.14
Qwen3-Omni 91.13 94.26 97.04 82.10
@@
Qwen2-Audio-Instruct 59.28 86.30 57.06 39.29 51.56 22.50 59.75
StepAudio 77.08 91.10 88.82 75.00 65.62 32.50 68.55
StepAudio2Mini 81.29 97.26 88.82 65.48 71.88 45.00 79.87
Mimo-Audio 84.01 94.52 88.82 72.62 78.12 52.50 85.53
@@
GLM4-Voice 65.00 (-17.20) 82.98 (-1.28) 47.50 (-32.50) 82.40 (0.20) 85.00 (-0.62) 82.67 (-2.66) 83.20 (3.20) 79.44 (0.00) 96.40 (0.60) 96.80 (0.80) 96.00 (0.40) 39.20 (-55.20)
Kimi-Audio 45.60 (-30.80) 65.96 (-7.23) 25.83 (-53.74) 56.20 (-20.60) 54.38 (-23.74) 77.33 (18.66) 73.60 (-7.20) 36.67 (-43.89) 93.80 (-1.60) 93.20 (-1.90) 94.40 (-1.20) 87.80 (-4.40)
Qwen2.5-Omni-7B 75.60 (-5.40) 82.55 (1.27) 68.75 (-12.08) 77.80 (-3.40) 79.38 (-1.87) 81.33 (5.33) 74.40 (-8.00) 77.22 (-5.56) 90.40 (0.20) 93.20 (3.20) 87.60 (-2.80) 89.20 (-1.60)
Baichuan-Audio-Chat 60.40 (-0.80) 61.70 (-0.43) 59.17 (-1.26) 61.60 (-0.60) 61.88 (-0.62) 62.67 (-5.33) 68.00 (3.20) 56.67 (-1.11) 77.20 (3.00) 83.60 (13.20) 70.80 (-7.20) 60.00 (-7.80)
Qwen2-Audio-7B-Instruct 76.40 (-5.80) 80.43 (-1.35) 72.50 (-10.00) 81.60 (0.80) 79.38 (0.63) 80.00 (-6.67) 83.20 (0.00) 83.33 (5.00) 93.20 (-1.40) 93.60 (-1.20) 92.80 (-1.60) 77.80 (-7.00)
StepAudio 62.53 (-18.94) 77.87 (-5.53) 47.50 (-32.08) 71.48 (-3.71) 68.75 (3.75) 76.00 (-13.33) 74.40 (-0.80) 70.00 (-8.33) 95.80 (1.00) 96.00 (1.20) 95.60 (0.80) 38.26 (-54.78)
StepAudio2 75.60 (-4.80) 75.32 (-2.13) 75.83 (-7.50) 75.40 (-2.00) 78.75 (-0.63) 88.00 (2.67) 67.20 (-4.80) 72.78 (-3.33) 96.60 (2.80) 96.00 (2.00) 97.20 (3.60) 76.00 (-14.40)
Mimo-Audio 76.00 (-14.80) 92.77 (-0.42) 59.58 (-28.75) 88.40 (-0.20) 90.62 (2.50) 96.00 (10.67) 88.80 (1.60) 82.78 (-8.33) 99.20 (1.00) 98.80 (-0.40) 99.60 (2.40) 60.40 (-33.00)
GPT4o-Audio 79.00 (-8.60) 88.94 (0.00) 69.17 (-17.08) 88.40 (2.20) 91.88 (4.38) 89.33 (0.00) 85.60 (0.80) 86.67 (2.23) 96.80 (-1.20) 97.20 (0.00) 96.40 (-2.40) 43.40 (-48.40)
Qwen3-Omni 85.47 (-9.48) 96.60 (1.71) 74.58 (-20.42) 95.19 (1.12) 93.12 (3.74) 98.67 (2.67) 96.80 (0.80) 94.44 (-1.67) 98.80 (-0.80) 98.80 (-0.80) 98.80 (-0.80) 95.22 (-3.91)
@@
GLM4-Voice 64.00 (-22.60) 45.00 (-36.75) 96.57 (-0.57) 74.86 (-12.57) 89.20 (-0.80) 85.56 (-3.89) 82.94 (-1.47) 93.03 (-0.68) 92.82 (1.57) 78.40 (-14.60)
Kimi-Audio 74.00 (-5.40) 60.50 (-11.65) 94.86 (4.57) 84.00 (-1.14) 75.80 (-9.60) 76.73 (-8.36) 58.24 (-11.47) 93.71 (0.45) 70.25 (-19.75) 76.80 (-11.20)
Qwen2.5-Omni-7B 69.40 (-16.80) 52.00 (-30.50) 95.43 (0.57) 82.86 (-2.85) 85.20 (-4.80) 80.74 (-6.17) 74.71 (-8.23) 95.51 (-1.79) 85.57 (-4.43) 79.80 (-10.20)
Baichuan-Audio-Chat 56.60 (-15.60) 45.00 (-20.00) 81.14 (-8.57) 58.29 (-12.57) 73.40 (-2.80) 73.45 (0.36) 64.41 (0.88) 81.35 (-5.39) 72.00 (-5.50) 67.20 (-4.60)
Qwen2-Audio-7B-Instruct 59.80 (-17.60) 37.75 (-32.00) 93.71 (-1.72) 76.00 (-1.14) 87.60 (-1.60) 82.18 (-6.18) 80.00 (-1.49) 96.18 (-0.18) 88.50 (0.25) 77.80 (-13.40)
StepAudio 55.87 (-30.80) 36.00 (-46.25) 93.71 (-0.58) 63.43 (-25.71) 86.92 (-2.94) 83.27 (-10.18) 77.65 (-1.47) 97.75 (0.45) 85.25 (-3.00) 66.50 (-27.25)
StepAudio2 65.00 (-19.20) 45.75 (-33.00) 94.86 (-1.14) 78.86 (-6.28) 83.40 (-5.80) 84.00 (-2.18) 66.18 (-12.06) 94.83 (-2.47) 84.75 (-6.75) 81.60 (-8.40)
Mimo-Audio 73.20 (-20.20) 54.00 (-35.50) 99.43 (4.43) 90.86 (-4.57) 90.40 (-5.00) 86.91 (-9.09) 86.76 (-5.00) 99.55 (0.45) 85.50 (-8.25) 95.20 (-4.20)
GPT4o-Audio 68.20 (-19.20) 61.00 (-20.50) 98.29 (-0.57) 54.86 (-34.28) 92.00 (-1.60) 86.18 (-6.55) 87.94 (-0.30) 96.40 (0.22) 94.25 (-1.50) 74.60 (-21.60)
Qwen3-Omni 83.47 (-13.86) 71.25 (-25.50) 97.71 (-1.72) 97.14 (0.57) 96.44 (-1.85) 96.73 (-1.45) 94.71 (-2.94) 100.00 (0.00) 93.75 (-3.25) 96.25 (-3.75)
@@
Avg. Casual Talk Mispronunciation Grammatical Error Topic Shift Code Switching
GLM4-Voice 84.41 (-6.90) 85.81 (-4.04) 82.20 (-10.60) 85.60 (-6.00) 86.20 (-6.40) 80.80 (-10.80)
Kimi-Audio 77.71 (-9.40) 81.18 (-4.44) 61.80 (-23.20) 77.40 (-10.40) 79.40 (-6.80) 84.00 (-8.80)
Qwen2.5-Omni-7B 80.44 (-6.75) 83.25 (-2.17) 80.60 (-8.40) 79.20 (-8.60) 75.00 (-11.40) 80.40 (-9.20)
Baichuan-Audio-Chat 67.46 (-7.02) 68.57 (-5.52) 71.40 (-1.20) 65.60 (-11.00) 61.60 (-12.40) 68.40 (-7.60)
Qwen2-Audio-7B-Instruct 79.60 (-9.04) 87.29 (-2.56) 78.40 (-10.00) 79.40 (-9.20) 67.00 (-19.80) 76.40 (-11.60)
StepAudio 85.81 (-5.64) 86.21 (-3.35) 89.44 (-2.69) 90.43 (-2.32) 82.42 (-7.78) 81.30 (-13.90)
StepAudio2 83.04 (-4.79) 85.02 (-2.27) 82.20 (-5.60) 80.80 (-8.20) 80.60 (-3.00) 83.60 (-8.60)
Mimo-Audio 95.10 (-3.04) 96.45 (-0.69) 96.00 (-3.60) 97.40 (-0.60) 92.00 (-5.60) 92.60 (-7.00)
GPT4o-Audio 92.34 (-1.94) 92.12 (-0.29) 92.60 (-4.00) 89.00 (-5.00) 90.60 (-2.60) 96.80 (-0.60)
Qwen3-Omni 97.61 (-1.62) 98.92 (-0.19) 94.61 (-4.04) 95.36 (-3.77) 96.92 (-2.42) 100.00 (0.00)